### v1

# MGAug: Multimodal Geometric Augmentation in Latent Spaces of Image Deformations

Abstract
Geometric transformations have been widely used to augment the size of training images. Existing methods often assume a unimodal distribution of the underlying transformations between images, which limits their power when data with multimodal distributions occur. In this paper, we propose a novel model, Multimodal Geometric Augmentation (MGAug), that for the first time generates augmenting transformations in a multimodal latent space of geometric deformations. To achieve this, we first develop a deep network that embeds the learning of latent geometric spaces of diffeomorphic transformations (a.k.a. diffeomorphisms) in a variational autoencoder (VAE). A mixture of multivariate Gaussians is formulated in the tangent space of diffeomorphisms and serves as a prior to approximate the hidden distribution of image transformations. We then augment the original training dataset by deforming images using randomly sampled transformations from the learned multimodal latent space of VAE. To validate the efficiency of our model, we jointly learn the augmentation strategy with two distinct domain-specific tasks: multi-class classification on 2D synthetic datasets and segmentation on real 3D brain magnetic resonance images (MRIs). We also compare MGAug with state-of-the-art transformation-based image augmentation algorithms. Experimental results show that our proposed approach outperforms all baselines by significantly improved prediction accuracy. Our code is publicly available at https://github.com/tonmoy-hossain/MGAug.
Introduction
Recent advances in deep learning (DL) have led to remarkable progress in image analysis tasks, including but not limited to image classification (Krizhevsky, Sutskever, and Hinton 2017; Wang and Zhang 2022), segmentation (Hossain et al. 2019; Chen et al. 2021), object detection (Kim and Pavlovic 2016; Pang et al. 2019), and 3D reconstruction (Jayakumar, Hossain, and Zhang 2023). The literature has shown that deep neural networks (DNNs) are most powerful when trained on vast amounts of manually labeled data. Unfortunately, many image applications have relatively small annotated datasets available, due to privacy concerns, regulations, costs, and incidents of disease. The DNNs, in such cases, tend to overfit due to their poor generalizations to large data variances. Data augmentation (DA) is a simple, yet efficient and commonly used strategy to overcome this issue by synthesizing more training examples to artificially increase the dataset (Cubuk et al. 2019; Zhao et al. 2019; Zoph et al. 2020; Shen et al. 2020). It is often employed as a preprocessing step applied to the training data to expose the DNNs to various types of approximated invariance; hence reducing the problem of overfitting. In the image domain, common augmentation strategies include adding variations to image intensities (Gao et al. 2021), or deforming the image by randomly generated geometric transformations (i.e., translation, rotation, or scaling) (Moreno-Barea et al. 2018). This paper focuses on DA, particularly in geometric spaces.
In practice, DA is usually a manual process where a human specifies a set of affine or deformable transformations believed to make DL models invariant (Hendrycks et al. 2019; Yun et al. 2019). However, such methods can pose a high risk of producing unrealistic or implausible variations of the image. For example, rotating an image 180 degrees makes sense for natural images, but produces ambiguities on digital numbers (i.e., the digits ‘6’ and ‘9’). In medical domain, randomly generated deformations may transform images from a healthy to a diseased group. To alleviate this issue, researchers have developed a group of automatic DA methods that learn an optimal augmenting policy from the data itself (Cubuk et al. 2019; Lim et al. 2019; Zoph et al. 2020; Araslanov and Roth 2021; Muralikrishnan et al. 2022). One approach is to learn a generator via generative adversarial networks (GANs) to synthesize data from scratch (Mirza and Osindero 2014; Chen et al. 2022). However, these models are difficult to train due to the high complexity of largely unknown parameter spaces (Jaderberg et al. 2015). Despite showing encouraging results on relatively easy datasets (e.g., $28^{2}$ MNIST), they do not appear to be applicable to more complex datasets with large structural and geometric variances. Another common approach is to learn a probabilistic distribution of a set of deformable transformations derived from the training images (Hauberg et al. 2016; Krebs et al. 2019; Dalca et al. 2019; Zhao et al. 2019; Olut et al. 2020; Shen et al. 2020). Most importantly, these methods assume a unimodal distribution of the underlying transformations between images. They do not provide an appropriate way to learn the full space of geometric transformations among the training data. This greatly limits their power when multimodal data (e.g., time-series/longitudinal images, or images of multiple sub-populations) occurs.
In this paper, we present a novel model, MGAug, that for the first time learns augmenting transformations in a multimodal latent space of geometric deformations. In contrast to existing approaches limited to data with unimodal distributions, our MGAug features the main advantages of
  1. (i)
    Generalizing the automatic DA process into a geometric space that allows hidden transformations with multiple modes.
  2. (ii)
    Developing a new generative model that learns the latent distribution of augmented transformations via a mixture of multivariate Gaussians defined as a prior in the tangent space of diffeomorphisms.
  3. (iii)
    Improving the quality of augmented samples from learned knowledge of training data itself with intra-group constraints.
We demonstrate the effectiveness of our proposed model in two distinct domain-specific tasks: multi-class classification and segmentation on both 2D synthetic data and real 3D brain magnetic resonance images (MRIs), respectively. We also compare MGAug with the state-of-the-art transformation-based augmentation algorithms (Olut et al. 2020; Dalca et al. 2019). Experimental results show that our method outperforms all baseline algorithms by significantly improved prediction accuracy.
Background: Geometric Deformations via Template-based Image Registration
We first briefly review the concept of template-based deformable image registration (Joshi et al. 2004; Zhang, Singh, and Fletcher 2013), which is widely used to estimate deformable geometric deformations from groupwise images. In this work, we adopt the algorithm of large deformation diffeomorphic metric mapping (LDDMM), which encourages the smoothness of deformation fields and preserves the topological structures of images (Beg et al. 2005).
Given a number of $N$ images $\{I_{1},\cdots,I_{N}\}$, the problem of template-based image registration is to find the shortest paths (a.k.a. geodesics) between a template image $I$ and the rest of image data with derived associate deformation fields $\phi_{1},\cdots,\phi_{N}$ that minimizes the energy function
$E(I,\phi_{n})=\sum_{n=1}^{N}\frac{1}{\sigma^{2}}\text{Dist}[I\circ\phi_{n}(v_{n}(t)),I_{n}]\,+\,\text{Reg}[\phi_{n}(v_{n}(t))],$ (1)
where $\sigma^{2}$ is a noise variance and $\circ$ denotes an interpolation operator that deforms image $I$ with an estimated transformation $\phi_{n}$. The $\text{Dist}[\cdot,\cdot]$ is a distance function that measures the dissimilarity between images, i.e., sum-of-squared differences (Beg et al. 2005), normalized cross correlation (Avants et al. 2008), and mutual information (Wells III et al. 1996). The ${\rm Reg}[\cdot]$ is a regularizer that guarantees the smoothness of transformations, i.e.,
[…]
$\text{with}\quad\frac{d\phi_{n}(t)}{dt}=-D\phi_{n}(t)\cdot v_{n}(t),$ (2)
where $L:V\rightarrow V^{*}$ is a symmetric, positive-definite differential operator that maps a tangent vector $v_{n}(t)\in V$ into its dual space as a momentum vector $m_{n}(t)\in V^{*}$. We write $m_{n}(t)=Lv_{n}(t)$, or $v_{n}(t)=Km_{n}(t)$, with $K$ being an inverse operator of $L$. The operator $D$ denotes a Jacobian matrix and $\cdot$ represents element-wise matrix multiplication. In this paper, we use a metric of the form $L=-\alpha\Delta+\mathbb{I}$, in which $\Delta$ is the discrete Laplacian operator, $\alpha$ is a positive regularity parameter, and $\mathbb{I}$ denotes an identity matrix.
The minimum of Eq. (2) is uniquely determined by solving a Euler-Poincaré differential equation (EPDiff) (Arnold 1966; Miller, Trouvé, and Younes 2006) with a given initial condition of velocity fields, noted as $v_{n}(0)$,
$\frac{\partial v_{n}(t)}{\partial t}$ $=-K[(Dv_{n}(t))^{T}\cdot m_{n}(t)+Dm_{n}(t)\cdot v_{n}(t)$
$+m_{n}(t)\cdot\operatorname{div}v_{n}(t)],$ (3)
where the operator $D$ denotes a Jacobian matrix, $\operatorname{div}$ is the divergence, and $\cdot$ represents an element-wise matrix multiplication. This is known as the geodesic shooting algorithm (Vialard et al. 2012), which nicely proves that the deformation-based shape descriptor, $\phi_{n}$, can be fully characterized by an initial velocity field, $v_{n}(0)$.
We are now able to equivalently minimize the energy function in Eq. (1) as
$E(I,v_{n}(0))$ $=\sum_{n=1}^{N}\frac{1}{\sigma^{2}}\text{Dist}[I\circ\phi_{n}(v_{n}(0)),I_{n}]$
[…]
For notation simplicity, we will drop the time index in the following sections.
Our Method: MGAug
This section presents a novel algorithm, MGAug, that learns a multimodal latent distribution of geometric transformations represented by the initial velocity fields $v_{0}$. We first introduce a generative model in latent geometric spaces for DA based on VAE, followed by an approximated variational inference. The augmented data is then used to train various image analysis tasks, i.e., image classification or segmentation. Similar to (Zhao et al. 2019; Chen et al. 2022), we employ a joint learning scheme to optimize the MGAug and IAT network parameters. To simplify the notation, we omit the time index of $v_{0}$ in the following sections.
Generative Model in Latent Multimodal Geometric Spaces
Given a group of $N$ input images $\mathbf{J}$ with a template $I$, we assume the latent geometric transformations $\mathbf{x}$ are generated from a number of $C$ clusters. The distribution of $\mathbf{x}$ can be formulated as a Gaussian mixture with means $\boldsymbol{\mu}$ and variances $\mathbf{\Sigma}$, specified by a VAE encoder parameterized by $\beta$, i.e., $p_{\theta}(\mathbf{x}|{\boldsymbol{\epsilon}})$, where $\boldsymbol{\epsilon}$ is a random number generated from $\mathcal{N}(0,\mathbb{I})$. That is to say, the VAE encoder outputs a set of $C$ means $\boldsymbol{\mu}$ and $C$ variances $\mathbf{\Sigma}$. A one-hot vector $\mathbf{z}$ is sampled from the mixing probability $\mathbf{\boldsymbol{\pi}}\in[0,1]$ with $\sum_{c=1}^{C}\mathbf{\boldsymbol{\pi}}^{c}=1$, which chooses one component from the Gaussian mixture. In this paper, we set the parameter $\mathbf{\boldsymbol{\pi}}=C^{-1}$ to make the variable $\mathbf{z}$ uniformly distributed, which is similar to (Dilokthanakul et al. 2016). The initial velocity field $\mathbf{v}$ is generated from a VAE decoder parameterized by $\theta$ and the continuous latent variable $\mathbf{x}$. The entire process to generate the observed image data $\mathbf{J}$ from a set of latent variables $\{\mathbf{v},\mathbf{x},\mathbf{\boldsymbol{\epsilon}},\mathbf{z}\}$ is:
$\mathbf{\boldsymbol{\epsilon}}$ $\sim\mathcal{N}(0,\mathbb{I}),$ (5)
$\mathbf{z}$ $\sim\text{Mult}(\mathbf{\boldsymbol{\pi}}),$ (6)
$\mathbf{x}|\mathbf{z},\mathbf{\boldsymbol{\epsilon}}$ $\sim\prod_{c=1}^{C}\mathcal{N}(\mathbf{\boldsymbol{\mu}}_{\mathbf{z}_{c}}(\mathbf{\boldsymbol{\epsilon}},\beta),\mathbf{\Sigma}_{\mathbf{z}_{c}}(\mathbf{\boldsymbol{\epsilon}},\beta))^{\mathbf{z}_{c}},$ (7)
[…]
$\mathbf{J}|\mathbf{v},I,\mathbf{\lambda}$ $\sim\mathcal{N}(\mathbf{J}\,|\,I\circ\mathbf{\phi}(\mathbf{v}),\boldsymbol{\lambda}^{2}\mathbb{I}),$ (9)
where $\text{Mult}(\cdot)$ denotes a multinomial distribution, $\mathbf{\boldsymbol{\lambda}}^{2}$ is a noise variance in the observed image space, and an identity matrix $\mathbb{I}$ maps a matrix to be diagonal. A graphical representation of our proposed generative model is shown in Fig. 1.
Figure 1: Graphical representation of our proposed generative model in multimodal latent space.
To guarantee the smoothness of the geometric transformations, we define a multivariate Gaussian distribution on the generated initial velocity fields $\mathbf{v}$ in Eq. (8) as
$p(\mathbf{v}|\mathbf{x})\propto\exp\left(-\frac{1}{2}(\mathbf{\Sigma}^{-1}\mathbf{v}_{\theta}(\mathbf{x}),\mathbf{v}_{\theta}(\mathbf{x}))\right).$ (10)
Similar to (Zhang and Fletcher 2013; Zhang, Shao, and Fletcher 2015), we define the covariance matrix $\boldsymbol{\Sigma}$ as an inverse operator of the differential operator, $L$, which is used in the registration regularity in Eq. (2).
Considering that our input images are measured on a discrete grid, we formulate our image noise model as an i.i.d. Gaussian at each image voxel, i.e.,
$p(\mathbf{J}|\mathbf{v},I,\mathbf{\boldsymbol{\lambda}})=\frac{1}{(2\pi)^{M/2}(\boldsymbol{\lambda})^{M}}\exp\left(-\frac{\|I\circ\mathbf{\phi}(\mathbf{v})-\mathbf{J}\|_{2}^{2}}{2\boldsymbol{\lambda}^{2}}\right),$ (11)
where $M$ is the total number of image voxels. Putting together Eq. (5) to Eq. (11), we arrive at a generative model
$p_{\beta,\theta}(\mathbf{J},\mathbf{v},\mathbf{x},\mathbf{\boldsymbol{\epsilon}},\mathbf{z})=p(\mathbf{J}|\mathbf{v},I,\mathbf{\boldsymbol{\lambda}})\,p(\mathbf{v}|\mathbf{x})\,p(\mathbf{x}|\mathbf{z},\mathbf{\boldsymbol{\epsilon}})\,p(\mathbf{z})\,p(\mathbf{\boldsymbol{\epsilon}})$ (12)
Variational Lower Bound
We will train the generative model in Eq. (12) with the variational inference objective, i.e. the log-evidence lower bound (ELBO). With an assumption that the mean-field variational family $q(\mathbf{v},\mathbf{x},\mathbf{\boldsymbol{\epsilon}},\mathbf{z}\,|\,\mathbf{J})$ is an approximate to the posterior, we are able to formulate the ELBO as
$\mathcal{L}_{\text{ELBO}}=\mathbb{E}\left[\log\frac{p_{\beta,\theta}(\mathbf{J},\mathbf{v},\mathbf{x},\mathbf{\boldsymbol{\epsilon}},\mathbf{z})}{q(\mathbf{v},\mathbf{x},\mathbf{\boldsymbol{\epsilon}},\mathbf{z}\,|\,\mathbf{J})}\right]\$ (13)
We factorize the mean-field variational family as $\mathbf{q}=\prod_{i=1}^{n}q(\mathbf{v}_{i}|\mathbf{J}_{i},I,\boldsymbol{\lambda})\,q_{\psi_{\mathbf{x}}}(\mathbf{x}_{i}|\mathbf{v}_{i},\mathbf{J}_{i})\,q_{\psi_{\mathbf{\boldsymbol{\epsilon}}}}(\mathbf{\boldsymbol{\epsilon}}_{i}|\mathbf{v}_{i})\,p_{\beta}(\mathbf{z}_{i}|\mathbf{x}_{i},\mathbf{\boldsymbol{\epsilon}}_{i}),$ where $i$ denotes the $i$-th image. Each variational factor is parameterized with the recognition networks $\psi_{\mathbf{x}}$ and $\psi_{\mathbf{\boldsymbol{\epsilon}}}$ that output the parameters of the variational Gaussian distributions. Note that we will drop the index $i$ in the following sections for the purpose of notation simplicity.
Figure 2: An overview of our model MGAug for image analysis tasks.
Using previously defined Eq. (10) and Eq. (11), the posterior $q(\mathbf{v}_{i}|\mathbf{J}_{i},I,\boldsymbol{\lambda})\propto p(\mathbf{J}|\mathbf{v},I,\mathbf{\boldsymbol{\lambda}})\cdot p(\mathbf{v}|\mathbf{x})$ is formulated by Baye’s rule. Following (Jiang et al. 2016), we can derive the posterior $p_{\beta}(\mathbf{z}|\mathbf{x},\mathbf{\boldsymbol{\epsilon}})$ as
$\centering p_{\beta}(\mathbf{z}_{j}=1|\mathbf{x},\mathbf{\boldsymbol{\epsilon}})\@add@centering$ $=\frac{p(\mathbf{z}_{j}=1)\,p(\mathbf{x}|\mathbf{z}_{j}=1,\boldsymbol{\epsilon})}{\sum_{c=1}^{C}p(\mathbf{z}_{c}=1)\,p(\mathbf{x}|\mathbf{z}_{j}=1,\boldsymbol{\epsilon})}$
$=\frac{\boldsymbol{\pi}_{j}\,\mathcal{N}(\mathbf{x}|\boldsymbol{\mu}_{j}(\boldsymbol{\epsilon};\beta),\boldsymbol{\Sigma}_{j}(\boldsymbol{\epsilon};\beta))}{\sum_{c=1}^{C}\boldsymbol{\pi}_{\mathbf{z}_{c}}\,\mathcal{N}(\mathbf{x}|\mu_{\mathbf{z}_{c}}(\boldsymbol{\epsilon};\beta),\Sigma_{\mathbf{z}_{c}}(\boldsymbol{\epsilon};\beta))},$
where $\mathbf{z}_{j},j\in\{1,\cdots,C\}$ denotes the $j$-th element of $\mathbf{z}$. The variational lower bound can be now formulated as
$\mathcal{L}_{ELBO}$ $=\mathbb{E}_{q(\mathbf{x}|\mathbf{J})}[\log P_{\theta}(\mathbf{J}|\mathbf{x})]$
$-\mathbb{E}_{q(\boldsymbol{\epsilon}|\mathbf{J})\,p(\mathbf{z}|\mathbf{x},\boldsymbol{\epsilon})}[KL(q_{\psi_{\mathbf{x}}}(\mathbf{x}|\mathbf{J})\,||\,p_{\beta}(\mathbf{x}|\boldsymbol{\epsilon},\mathbf{z}))]$
$-KL(q_{\psi_{\mathbf{\boldsymbol{\epsilon}}}}(\boldsymbol{\epsilon}|\mathbf{J})\,||\,p(\boldsymbol{\epsilon}))$
$-\mathbb{E}_{q(\mathbf{x}|\mathbf{J})\,q(\boldsymbol{\epsilon}|\mathbf{J})}[KL(p_{\beta}(\mathbf{z}|\mathbf{x},\boldsymbol{\epsilon})\,||\,p(\mathbf{z}))]$ (14)
Joint Learning of MAGug With Image Analysis
We demonstrate the effectiveness of MGAug in image analysis tasks (IAT) by developing a joint learning framework. It has been shown in (Zhao et al. 2019; Chen et al. 2022) that these two tasks can be mutually beneficial. The MGAug effectively synthesizes image samples to augment the training dataset, while the IAT can provide extra information on rectifying the learned distributions of latent geometric spaces. An overview of our proposed network architecture is shown in Fig. 2.
Network design of MGAug.
MGAug is mainly designed based on the previously introduced inference model, where the approximate posteriors are simulated by neural networks. Analogous to (Kingma and Welling 2013), we let the encoder of MGAug, parameterized by $\beta$, approximate the posterior as a mixture of multivariate Gaussians. In the latent space $\mathbf{z}$, we use non-linear hidden layers to learn the parameters (i.e., means and covariances) of the mixture Gaussians. The decoder, parameterized by $\theta$, outputs augmented velocity fields $\mathbf{v}_{n}$, followed by generating their associated deformations $\phi_{n}$ (via Eq. (3)) and augmented deformed images. In this paper, we adopt UNet architecture (Ronneberger, Fischer, and Brox 2015) as the network backbone for MGAug. However, other networks such as UNet++ (Zhou et al. 2018) and TransUNet (Chen et al. 2021) can be easily applied.
Defining $\Theta=(\mathbf{J},\beta,\theta,\mathbf{x},\mathbf{z},\mathbf{v},\boldsymbol{\epsilon},I)$ as posterior parameters, we formulate the network loss of MGAug as a combination of dissimilarity between deformed images and templates, a regularity to ensure the smoothness of augmenting transformations, and the variational inference objective as
$\ell_{\text{MGAug}}(\Theta)$ $=-\log p_{\beta,\theta}(\mathbf{J},\mathbf{v},\mathbf{x},\mathbf{\boldsymbol{\epsilon}},\mathbf{z})$
$\approx\frac{\|I\circ\mathbf{\phi}(\mathbf{v}(\beta,\theta))-\mathbf{J}\|_{2}^{2}}{2\boldsymbol{\lambda}^{2}}+\mathcal{L}_{ELBO}$
$+\frac{1}{2}(L\mathbf{v}(\beta,\theta),\mathbf{v}(\beta,\theta))+\text{reg}(\beta,\theta),$
where $\text{reg}(\cdot)$ denotes network parameter regularizations.
Network design of IAT.
We demonstrate MGAug on the task of image classification and segmentation separately. Other domain-specific tasks, such as image reconstruction, object recognition, and tracking can be integrated similarly.
For classification, assuming we have augmented images (consist of $C$ classes with $N_{c}$ images per class) and their labels denoted as $\hat{y}$. The classifier is trained by minimizing the loss
$\ell_{\text{IAT}}^{clf}=-\gamma\sum_{n=1}^{N_{c}}\sum_{c=1}^{C}\hat{y}_{nc}\cdot\log[y_{nc}(\beta,\theta,w_{1})]+\text{reg}(\beta,\theta,w_{1}),$ (15)
where $y_{nc}$ denotes the predicted label, $\gamma$ is the weighting parameter, and $w_{1}$ is a classification network parameter.
For segmentation, we employ a loss function as a combination of Sørensen--Dice coefficient (Dice 1945) and cross-entropy (CE). Let $Q$ denote the number of segmentation labels, $S_{q}$ denotes the augmented ground truth, and $R_{q}$ denotes the prediction with the associated probabilistic map $R^{p}_{q}$ for $q$-th label. We formulate the segmentation loss as
$\ell_{\text{IAT}}^{seg}$ $=\tau\sum_{q=1}^{Q}[1-\text{Dice}(S_{q},R_{q}(\beta,\theta,w_{2}))$
$+\text{CE}(S_{q},R^{p}_{q}(\beta,\theta,w_{2}))]+\text{reg}(\beta,\theta,w_{2}),$ (16)
where $\text{Dice}(S_{q},R_{q})=2(|S_{q}|\cap|R_{q}|)/(|S_{q}|+|R_{q}|)$ and $\text{CE}(S_{q},R^{p}_{q})=-S_{q}\cdot\log R_{q}$. The $\tau$ is the weighting parameter and $w_{2}$ denotes the segmentation network parameters.
The total loss of our joint learning framework of MGAug and IAT is $\ell=\ell_{\text{MGAug}}+\ell_{\text{IAT}}$. We employ an alternative optimization scheme (Nocedal and Wright 1999) to minimize the loss by alternating between the training of the MGAug network and the IAT network without sharing weights, making it an end-to-end network architecture (Zhao et al. 2019; Chen et al. 2022).
Experimental Evaluation
We validate our proposed model, MGAug, on both 2D synthetic datasets and real 3D brain MRI scans, highlighting its effectiveness on training data with limited availability.
Datasets
2D silhouette shapes. We choose Kimia216 (Sebastian, Klein, and Kimia 2004), which is a relatively smaller dataset with seven classes of 2D silhouette shape data (bone, bottle, children, face, fish, heart, and key). Each shape class includes approximately $20$ binary images, with the size of $28^{2}$. The total number of our training images for multi-classification is about $140$.
2D handwritten digits. We use $70000$ handwritten digits of $10$ classes (00-$9$) from the MNIST repository (LeCun 1998). Each class includes approximately $7000$ images with the size of $28^{2}$.
3D brain MRIs. We include $803$ subjects of public T1-weighted longitudinal brain MRIs with manually delineated anatomical structures from Open Access Series of Imaging Studies (OASIS) (LaMontagne et al. 2019), Alzheimer’s Disease Neuroimaging Initiative (ADNI) (Jack Jr et al. 2008), Autism Brain Imaging Data Exchange (ABIDE) (Di Martino et al. 2014), and LONI Probabilistic Brain Atlas Individual Subject Data (LPBA40) (Shattuck et al. 2008). Due to the difficulty of preserving the diffeomorphic property across individual subjects with large age variations, we carefully evaluate images from subjects aged from $60$ to $90$. All MRIs were pre-processed as $256\times 256\times 256$, $1.25mm^{3}$ isotropic voxels, underwent skull-stripping, intensity normalization, bias field correction and pre-alignment with affine transformations.
Experiments and Implementation Details
We validate the proposed model, MGAug, on both image classification and segmentation tasks. As MGAug falls into the category of transformation-based DA schemes, we mainly focuses on comparing its performance with similar baselines: randomly generated affine transformations (ATA, a commonly used augmentation strategy), and two state-of-the-art unimodal generative models based on diffeomorphic registrations (AdvEigAug or AEA, an augmentation model based on the statistics of a learned latent space of velocity fields (Olut et al. 2020); and UDA, a unimodal deformation-based augmentation using random stationary velocity fields (Dalca et al. 2019)).
Evaluation of classification on small datasets. We first validate our model and other baselines on a relatively small dataset of 7-class 2D silhouette shapes. We adopt six classification network backbones, including a three-layer CNN, ResNet18 (He et al. 2016), VGGNet16 (Simonyan and Zisserman 2014), SqueezeNet (Iandola et al. 2016), DenseNet121 (Huang et al. 2017), and Vision Transformer (ViT) (Dosovitskiy et al. 2020).
We carefully determine the optimal number of modes, $C$, for our multimodal generative model by performing cross-validation with various values of $C$. The value of C will vary for different image analysis tasks depending on the dataset distributions. For classification, we assume that the value of C is close to the number of classes. Furthermore, we conduct a comprehensive assessment of the augmentation efficiency offered by our proposed approach. This involves augmenting the training dataset to various magnitudes, ranging from $2$ to $10$ times the size of the original dataset.
To examine the quality of the generated augmenting transformations, we compute the determinant of Jacobian of the deformation fields. Throughout the evaluation process, we employ multiple metrics such as average accuracy (Acc), precision (Prec), and F1-score (F1) to validate the classification performance.
Evaluation of segmentation on real 3D brain MRIs. We validate the performance of all methods on real 3D brain images with a decent size of training data. In this experiment, we employ TransUNet as our network backbone, given its proven superiority in brain segmentation performance (Chen et al. 2021). To ensure a fair comparison of the final segmentation accuracy, we augment the training data three times for all methods, resulting in a total number of approximately $2000$ augmented images. We then compare the segmentation dice scores of all augmentation methods over different brain structures.
Evaluation of the quality of augmented images. To further evaluate the quality of the augmented images, we validate MGAug on two relatively large datasets (2D MNIST and 3D brain MRIs) with various levels of reduced training sample sizes (i.e., $10\%$, $20\%$, and $50\%$ of the original size). We augment the training images three times of the original size for all methods.
Parameter setting. We set parameter $\alpha=3.0$ for the operator $L$ and the number of time steps for Euler integration in EPDiff (Eq. (3)) as $10$. We set the noise variance $\sigma$ as $0.02$, the batch size as $16$, and use the cosine annealing learning rate scheduler that starts from a learning rate $\eta=1e^{-3}$ for network training. We run all the models on training and validation images with Adam optimizer until obtaining the best validation performance. The training and prediction procedure of all learning-based methods are performed on a 40GB NVIDIA A100 Tensor Core GPU. For both 2D and 3D datasets, we split the images by using $70\%$ as training, $15\%$ as validation, and $15\%$ as testing images.
Our method is flexible to either automatically estimated or pre-selected template(s) for group-wise data. Therefore, we demonstrate estimated with-in-class template(s) for classification tasks vs. pre-selected templates for segmentation tasks, respectively. In all experiments presented in this paper, we run cross-validation to determine the optimal number of modes $C$.
Results
The left panel of Fig. 3 reports our model MGAug’s classification accuracy (using ViT as the network backbone) on 2D silhouette shape dataset with different number of modes, $C$, along with diverse levels of augmented images quantities (i.e., 5×,7×,5$\times$,7$\times$,5 × , 7 × , and 10×10$\times$10 × the original training data size). It suggests that the utilization of a multimodal latent space $(C>1)$ consistently yields superior performance compared to an unimodal distribution $(C=1)$ by more than $2$-$7\%$. Our model achieves an optimal performance when $C=7$, which represents the number of classes in the training data. The right panel of Fig. 3 shows a comprehensive evaluation of classification accuracy across all augmentation methods with an increased number of augmented images in training data. Our method significantly outperforms other baselines by more than $7\%$ of accuracy with different numbers of augmented images. Another interesting observation is that our MGAug achieves comparable classification accuracy when using roughly 3×3$\times$3 × augmented images for training vs. other baselines using 10×10$\times$10 × of ground truth images. This indicates the potency of our method in efficiently leveraging augmented data to achieve competitive results.
Table 1: Accuracy comparison on 2D shapes over all models with various network backbones.
ATA AEA UDA MGAug
Acc Prec F1 Acc Prec F1 Acc Prec F1 Acc Prec F1
CNN 71.73 73.25 66.89 83.93 88.82 83.07 85.71 87.14 83.73 91.07 92.14 90.78
ResNet18 76.78 79.22 73.44 80.36 82.17 78.76 87.50 90.07 86.83 89.28 90.43 89.15
VGGNet16 73.21 72.46 71.23 85.71 88.04 85.85 85.71 88.57 81.58 89.28 91.21 89.22
SqueezeNet 75.00 75.85 72.99 76.78 79.90 74.53 83.93 87.30 82.73 92.86 93.97 92.59
DenseNet121 75.89 77.54 73.88 87.50 90.79 86.18 89.28 90.86 89.08 94.64 95.55 94.28
ViT 78.57 82.62 76.53 89.28 92.20 89.33 91.07 91.93 96.84 98.21 98.41 98.20
Figure 3: Left: Accuracy comparison on various modes $C$ on 2D shape dataset. Right: A comparison of classification performance for all models over increasing number of augmented data taking all ground truth images.
Tab. 1 reports the multi-class classification accuracy on a relatively smaller 2D silhouette shape data across all augmentation methods with various classification network backbones. For a fair comparison, we augment the data 10×10$\times$10 × of the original training images. MGAug achieves the highest performance, outperforming all baselines by $2$-$9\%$ of accuracy on six different network backbones.
Fig. 4 visualizes examples of augmented images and the determinant of Jacobian (DetJac) of augmented transformations produced by MGAug. The value of DetJac indicates volume changes when applying augmented deformations on templates. For example, there is no volume change when DetJac$=$1, while volume shrinks when DetJac$<$1 and expands when DetJac$>$1. The value of DetJac smaller than zero indicates an artifact or singularity in the transformation field, i.e., a failure to preserve the diffeomorphic property when the effect of folding and crossing grids occurs. Our method nicely preserves the topological structures of augmented images without introducing artifacts.
Figure 4: Examples of DetJac generated by MGAug. Left to right: source/template, target images, augmented DetJac overlaid with template images.
Fig. 5 reports the statistics of dice scores (mean and variance) of all augmentation methods over fifteen brain structures experimenting on $30\%$ of training images with $2000$ augmented images. Our MGAug outperforms all the baselines with a significant dice improvement in all anatomical structures (at least $2\%$ or higher).
Figure 5: Dice score comparison over fifteen anatomical brain structures, (LWM - Left White Matter, RWM - Right White Matter, LLV - Left Lateral Ventricle, RLV - Right Lateral Ventricle, LCC - Left Cerebellum Cortex, RCC - Right Cerebellum Cortex, LT - Left Thalamus, RT - Right Thalamus, LC - Left Cudate, RC - Right Caudate, LP - Left Putamen, RP - Right Putamen, LH - Left Hippocampus, RH - Right Hippocampus, BS - Brain Stem) for four augmentation methods.
The left panel of Fig. 6 showcases the classification results on the 2D MNIST dataset for all methods, using different percentages of ground truth data during the training phase. Our MGAug method consistently achieves the highest accuracy across all augmentation techniques. Remarkably, MGAug reaches an accuracy of $98.02\%$ (which is the baseline accuracy without any augmentation) using only $20\%$ of the ground truth data, whereas other methods require at least $50\%$ of ground truth data to achieve similar results. Similarly, the right panel of Fig. 6 presents the dice evaluation for 3D brain MRI segmentations, as the percentage of ground truth data used in training is progressively increased. In a manner mirroring the classification results, our MGAug consistently outperforms the baseline algorithms by yielding the highest dice scores. These results indicate that our augmentation method has great potential to better alleviate the heavy need for labor and time-consuming ground-truths.
Figure 6: Left to right: classification evaluation on 2D MNIST dataset; segmentation evaluation on 3D brain MRIs with various percentages of ground truth used in the training inference. The horizontal dashed line (- -) serves as a reference, representing the performance achieved when utilizing all available training images without any augmentation.
Additional discussion.
It is to our attention that diffusion models have recently become a powerful generative model (Ho, Jain, and Abbeel 2020). In this context, we initiate a preliminary comparison between our MGAug model and data augmentation through a diffusion model operating in the image space (by manipulating image intensities and textures) (Trabucco et al. 2023). This comparison was carried out with a specific focus on classification tasks using the 2D MNIST dataset.
Experimental results show that the diffusion model achieves an accuracy of $97.56\%$ when using $50\%$ of the ground truth labels. This accuracy, interestingly, is slightly lower than what our MGAug method accomplishes ($98.02\%$) when only using $20\%$ of the ground truth data. This may indicate the advantages of multimodal distributions of latent geometric space developed in our model. Our future work will further explore the power of diffusion models in the latent geometric space for efficient data augmentation, which has not been investigated so far.
MGAug vs. intensity-based augmentations. While the primary emphasis of this paper does not fall into the realm of intensity-based augmentations, we compare MGAug with a few notable DA schemes that manipulate image intensities or textures. These methods include AutoAugment (Cubuk et al. 2019), RandAugment (Cubuk et al. 2020), AugMix (Hendrycks et al. 2019), and TrivialAugment (Müller and Hutter 2021).
Tab. 2 reports the accuracy comparison of a variety of DA schemes for image classification networks based on the ViT backbone. The testing results on both MNIST digits and 2D shape data show that our model consistently outperforms the baselines.
Table 2: Accuracy comparison on both 2D digits and 2D shape datasets over a mix of augmentation methods (including intensity-based vs. transformation-based DA) on ViT.
Datasets 2D Digits 2D Shape
Models Acc F1 Acc F1
AutoAugment 90.82 90.73 76.05 75.25
RandAugment 95.25 95.22 81.25 81.70
AugMix 97.31 97.30 74.28 69.35
TrivialAug 95.48 95.49 83.93 82.75
Affine 95.90 95.84 78.57 76.53
AEA 96.95 96.91 89.28 89.33
UDA 97.85 97.83 91.07 96.84
MGAug 98.62 98.78 98.21 98.20
Conclusion & Discussion
In this paper, we present a novel data augmentation model, MGAug, which learns augmenting transformations in a multimodal latent space of geometric deformations reparameterized by low-dimensional tangent spaces of diffeomorphisms. In contrast to existing augmentation models that adopt a unimodal distribution to augment data, our augmentation method fully characterizes the multimodal distribution of image transformations via introducing a mixture of multivariate Gaussians as the prior. Our method significantly improves the performance of multi-class classification (on 2D dataset) and image segmentation (on 3D brain MRIs) and eliminates the artifacts of geometric transformations by preserving topological structures of augmented images. While MGAug focuses on LDDMM, the theoretical tool developed in our model is widely applicable to various registration frameworks, for example, stationary velocity fields (Modat et al. 2012).
To the best of our knowledge, we are the first to learn geometric deformations in a multimodal manner via deep neural networks. Our augmentation model can be easily applied to different domain-specific tasks, e.g., image reconstruction and object tracking. Interesting future directions based on this work can be i) investigating the model interpretability for classification tasks when geometric deformations are learned in a multimodal manner, and ii) the robustness study of MGAug when image intensity variations occur, i.e., under adversarial noise attacks.
Acknowledgments.
This work was supported by NSF CAREER Grant 2239977.